OpenStudy (anonymous):

I have to memorize all the square numbers from 1-30 1^2 2^2 3^2 all the way to 30^2 Is there a trick to memorizing it or being able to solve it quick?

OpenStudy (anonymous):

Multiplying them... just that. No need to memorize them, you can do it by hand

OpenStudy (anonymous):

but my math teacher told me so.

OpenStudy (♪chibiterasu):

You could use flash cards.

OpenStudy (anonymous):

how can i do that in just 1 day ?

OpenStudy (anonymous):

write them over and over until you can memorize them all without looking at anything...get your parents to do flash cards with you or either get them to write them down for you and then time you for 1 minute to see how many you get correct :) good luck!!!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

but any other (easy) way ?

OpenStudy (freckles):

Well I remember in elementary school we had to re-memorize the alphabet and the time tables. Just use the same technique you chose to remember those to remember these.

OpenStudy (anonymous):

i don't have much time.

jimthompson5910 (jim_thompson5910):

Memorizing 1^2 through 10^2 shouldn't be that difficult (just practice a lot with them) To get 11^2, you could multiply 11*11 = 121 or you can add on 2x+1 = 2*10+1 = 21 to 100 to get 121 (notice how x = 10) now you know 11^2 = 121 ----------------- To get 12^2, you can multiply 12*12 = 144 or you can add on 2x+1 = 2*11+1 = 23 (now x = 11) ------------------ keep the pattern going 12^2 = 144 13^2 = 144+25 = 169 notice how 25 = 2*12+1 and how we have that 12 in there so basically you are adding odd numbers to the last perfect square to get the next perfect square 14^2 = 169+27 = 196 15^2 = 196+29 = 225 etc etc

OpenStudy (freckles):

You know if you don't actually won't to do the math each time.

OpenStudy (freckles):

also the squares from 1 to 30 shouldn't that just be 1^2 to 5^2?

OpenStudy (freckles):

1^2=1 2^2=4 3^2=9 4^2=16 5^2=25 because 6^2=36>30

OpenStudy (anonymous):

ahh ! but my teacher told me to do that.

OpenStudy (anonymous):

ok guys , well thanks for ur replies.